Based on the Consumer Price Index, the rate of inflation in the United States between 1980 and 2011 was about 3.3% per year. What salary in 1980 would be equivalent to a $51,000 salary in 2011? Round your final answer to two decimal places.
1
Expert's answer
2020-09-20T17:53:52-0400

solution


number of years (n) between 1980 and 2011



"n=2011-1980=31years"

Inflation rate (r) "=3.3\%" per year


Let X be the salary in 1980. The salary equivalent in 2011 is given by:



"X*(1+r)^n"

Therefore



"X*(1+0.033)^{31}= 51,000"


"X = \frac{51,000}{(1+0.033)^{31}}"

"=18,640.61"

answer: 1980 salary would be $18,640.61